IO Two polarizers reduce to intensity incident unpolarized light to crnly angle both een Z polarizers
Answer #1

Since the intial state of polarization is not give , I assume that the light is intially unpolarized hence, the intensity of light comming out of first polarizer is clearly\frac{I_o}{2} , if \theta is the angle between the polarizers ,using malus law we have

\frac{I_o}{10}=\frac{I_o}{2}\cos^2(\theta)\Rightarrow \cos\theta=\frac{1}{\sqrt{5}}\Rightarrow \theta=45^o